A ball with a mass 0.5 kg is dropped from a height of 10 m to the ground and bounces back to a height of 7m. How much potential energy does the ball lose


1
Expert's answer
2022-02-09T06:28:55-0500

Given:

"m=\rm 0.5 \: kg"

"h_1=\rm 10 \: m"

"h_2=\rm 7 \: m"


The loss in potential energy of the ball

"\Delta E_p=mg(h_1-h_2)\\
\rm=0.5\: kg*9.8\: N/kg*(10-7)\: m=14.7\: J"